# VideoChat-Flash: Hierarchical Compression for Long-Context Video Modeling

### v2

@@ 3.4 Multi-hop Needle In A Video Haystack
Figure 5: An example of “multi-hop needle in a video haystack”. Frames 1 to 3 and needle are the inserted indicating images while X1, X2, and X3 and decoy pictures for distraction. MLLMs are asked to both find the needle (by choosing its right description) and then answer its related question.
Needle in a video haystack requires the model to locate an indicated image in a long video, find a target word, or answer a question based on that image. Most existing benchmarks [75, 80] typically inject a target image into a long video, making it difficult to assess genuine long-context understanding. Models might succeed by simply finding the most visually distinct frame, instead of targeted searching according to instructions. This design results in quite high accuracy (often exceeding 99%) by almost every model on these benchmarks.
Regarding this, we propose a new benchmark, termed multi-hop needle in a video haystack. It rigorously evaluates the model’s ability to process extremely long contexts by demanding it to locate a sequence of interconnected indicating images embedded within a long video, as illustrated in Fig. 5. The first indicating image could be located via the given instruction, while subsequent images are found using clues embedded within the preceding images, creating a chain. Only after successfully navigating this whole chain and identifying the final “needle” image can the model answer a question related to that image.
To avoid trivial solutions, we give two constraints to ensure the answer can only be obtained through the designed order. First, we incorporate decoy images (three in our experiments) with misleading clues, forcing the model to follow the correct chain. Second, the model must both find the needle (by choosing the right option) and answer the corresponding question correctly, minimizing the probability of success by guessing. In practice, all images are sourced from MS-COCO [36], leveraging its human-annotated captions and question-answer pairs. This multi-hop approach provides a far more robust evaluation of long context understanding in multimodal large language models (MLLMs) than the current single-hop.
@@ 4 Experiments
In dynamic video sampling, we randomly sample 64-512 frames. For VideoChat-Flash, we employ UMT-L@224 [30], token merging with MLP, and Qwen2-7B as visual encoder, connector, and LLM, respectively. When processing a long video, we divide it into shorter clips, each consisting of 4 frames. Each clip is then compressed into 64 tokens, meaning that, on average, each frame is represented by 16 tokens. Regarding video-level compression, while it presents some challenges in compatibility with training acceleration strategies such as sequence parallelism (although our experiments confirm that enabling it during training does not compromise performance), we employ different configurations during inference depending on the use case. For efficiency-focused scenarios, we utilize higher drop rates and more aggressive drop methods, whereas for performance-oriented scenarios, we apply dropout only in the later layers of the LLM.
Model FLOPs Memory(G)
(T) Train Infer
Input 64 frames
LongVILA [66] 455.1 56.7 18.1
LongVA [75] 316.8 50.6 17.4
VideoChat-Flash (ours) 29.6 35.2 15.5
Input 256 frames
LongVILA [66] 2946.3 oom 26.8
LongVA [75] 1870.1 oom 23.7
VideoChat-Flash (ours) 113.3 40.2 16.1
Input 1000 frames
LongVILA [66] 28707.7 oom 79.7
LongVA [75] 16583.7 oom 48.3
VideoChat-Flash (ours) 519.7 59.4 18.5
Input 10000 frames
LongVILA [66] 2368810.0 oom oom
LongVA [75] 1289490.0 oom oom
VideoChat-Flash (ours) 15786.0 oom 47.4
Table 1: Comparison of FLOPs and Cuda memory. FLOPs and memory are estimated by the context length of the input LLM.
@@ 4.1 General Video Understanding Evaluation
Model Size #tokens MVBench PerceptionTest LongVideoBench MLVU VideoMME (w/o & w sub.) LVBench per frame Avg Val Val M-Avg Overall Long Avg Avg. Duration 16s 23s 473s 651s 1010s 2386s 4101s Proprietary Models GPT4-V [43] - - 43.7 - 59.1 49.2 59.9/63.3 53.5/56.9 - GPT4-o [44] - - 64.6 - 66.7 64.6 71.9/77.2 65.3/72.1 30.8 Gemini-1.5-Pro [47] - - 60.5 - 64.0 - 75.0/81.3 67.4/77.4 33.1 Small Size MLLMs Qwen2-VL [55] 2B - 63.2 - - - 55.6/60.4 - - InternVL2.5 [11] 2B 256 68.8 - 46.0 61.4 51.9/54.1 - - VideoChat-Flash @448 2B 16 70.0 70.5 58.3 65.7 57.0/63.9 44.9/54.0 42.9 Open-Source MLLMs VideoLLaMA2 [13] 7B 72 54.6 51.4 - 48.5 47.9/50.3 42.1/43.8 - VideoLLaMA2 [13] 72B 72 62.0 57.5 - - 62.4/64.7 57.6/59.0 - VideoChat2-HD [31] 7B 72 62.3 - - 47.9 45.3/55.7 39.8/53.9 - InternVideo2-HD [59] 7B 72 67.2 63.4 - - 49.4/ - - - IXComposer-2.5 [74] 7B 400 69.1 34.4 - 37.3 55.8/58.8 - - InternVL2 [12] 8B 256 65.8 - 54.6 64.0 54.0/56.9 - - InternVL2 [12] 76B 256 69.6 - 61.1 69.9 61.2/62.8 - - InternVL2.5 [11] 8B 256 72.0 - 60.0 68.9 64.2/66.9 - - Qwen2-VL [55] 7B - 67.0 62.3 - - 63.3/69.0 - - Qwen2-VL [55] 72B - 73.6 68.0 - - 71.2/77.8 - 41.8 LLaVA-NeXT-Video [78] 7B 144 53.1 48.8 49.1 - - /46.5 - - LLaVA-OneVision [27] 7B 196 56.7 57.1 56.3 64.7 58.2/61.5 - - LLaVA-OneVision [27] 72B 196 59.4 66.9 61.3 68.0 66.2/69.5 - - LLaVA-Video [79] 7B 676 58.6 67.9 58.2 70.8 63.3/69.7 - - Open-Source Long Video MLLMs LLaMA-VID [33] 7B 2 41.9 44.6 - 33.2 25.9/ - - 23.9 Kangaroo [38] 8B 256 61.0 - 54.8 61.0 56.0 / 57.6 46.7 / 49.3 39.4 LongVILA [66] 7B 196 67.1 58.1 57.1 - 60.1/65.6 47.0/52.1 - LongVA [75] 7B 144 - - - 56.3 52.6/54.3 46.2/47.6 - LongLLaVA [57] 9B 144 49.1 - - - 43.7/ - 36.4/ - - LongVU [50] 7B 64 66.9 - - 65.4 - /60.6 - /59.5 - VideoChat-Flash @224 7B 16 73.2 75.6 64.2 74.5 64.0/69.4 53.6/61.9 47.2 VideoChat-Flash @448 7B 16 74.0 76.2 64.7 74.6 65.3/69.7 55.4/63.3 48.2
Table 2: Results on comprehensive video question answer benchmarks
@@
We evaluate our model on five general video understanding benchmarks in question-answering format, including two short video benchmarks: MVBench [31] and Perception Test [45], and three long video benchmarks: LongVideoBench [64], MLVU [81] and LVBench [56], and a comprehensive benchmark, VideoMME [16], covering videos ranging from minute-level to hour-level durations. We further evaluate the model’s timestamp awareness using the Charades-STA temporal grounding dataset [17].
Short video understanding.
As shown in Tab. 2, our VideoChat-Flash gives the best results on MVBench (74.0%) and Perception Test (76.2%) among all, surpassing other approaches by a notable margin. We have achieved new state-of-the-art (SOTA) results among 7B models. Moreover, our performance even surpasses that of models with much larger scales, such as InternVL2-76B and VideoLLaMA2-72B, and proprietary models like GPT4o and Gemini-1.5-Pro. Furthermore, our model achieved a comprehensive lead in results at the 2B parameter scale, even surpassing some models with 7B or even 72B parameters. This demonstrates the effectiveness of our approach in enhancing video perception capabilities.
Long video understanding.
Tab. 2 gives that VideoChat-Flash not only outperforms previous open-source general MLLMs of the same model scale but also surpasses a host of MLLMs designed specifically for long videos on the long video understanding benchmarks. Especially, it gets the new SOTA performance (74.7%) on MLVU, even compared with GPT4-V and GPT4-o. This empirically proves that VideoChat-Flash with HiCo can handle long multimodal contexts effectively. More importantly, despite performance, the training and inference efficiency of our model is also much higher than that of previous ones. As in Tab. 1, even when processing short videos, the compute load of our model is only one-tenth that of previous models. Meanwhile, as the number of input frames increases, the difference becomes more and more pronounced. On a single A100 device, only our model can complete the inference on 10,000 frames (approximately equivalent to a 3-hour video). Concretely, VideoChat-Flash’s compute load is two orders of magnitude lower than that of LongVILA [66] (15,786 vs. 2,368,810 TFLOPs).
Method Charades-STA
$R@1_{(IOU=0.3)}$ $R@1_{(IOU=0.5)}$ $R@1_{(IOU=0.7)}$
Specialized Models
2D-TAN [77] 57.3 45.8 27.9
MMN [61] 65.4 53.3 31.5
MDETR [25] 65.8 52.1 30.6
UniVTG [35] 72.6 60.2 38.6
Open-Source MLLMs (7B size)
VTimeLLM [20] 51.0 27.5 11.4
HawkEye [60] 50.6 31.4 14.5
TimeChat [48] - 32.2 13.4
TimeSuite [69] 69.9 48.7 24.0
VideoChat2 [31] 9.6 3.4 1.4
LLaVA-OneVison [27] 31.2 13.5 5.2
VideoChat-Flash@224 74.5 53.1 27.6
VideoChat-Flash@448 72.5 51.4 26.4
Table 3: Results on temporal grounding task. Similar to previous Video MLLM approaches [48, 69], our model is trained without any grounding data from the Charades-STA dataset, evaluating its zero-shot capabilities.
Temporal Grounding
As detailed in Tab. 3, in contrast to previous approaches [48, 69] that rely on intricate architectures for timestamps awareness, our model exhibits surprisingly strong temporal grounding capabilities. Its zero-shot performance substantially exceeds that of existing MLLMs, even reaching or surpassing the performance of dedicated, fully supervised models.
4.2 Long Context Evaluation with NIAH
@@
Fig. 6 presents the results of baselines and ours on typical “single-hop Needle-in-a-haystack” for long videos, widely used in long video evaluations [75, 66, 57, 50]. Adhering to the protocols in LongVA [75], we source a long video and sample frames uniformly from it. Then we add needles (indicating images) into the sampled image sequence at different positions. MLLMs are fed with this long image sequence and answer the corresponding questions to the indicating images. Different from existing video NIAH with thousands of frames, we evaluate all models over 10,000 frames. Note our VideoChat-Flash delivers a 99.1% success rate in accurately retrieving the correct indicating image and answering the related question even across 10,000 frames. In comparison, LongVA gives a decent result close to 92% within 3000 frames while LLama-VID only achieves 55% accuracy. It demonstrates VideoChat-Flash’s state-of-the-art performance in long multimodal context modeling.
(a) LongVA [75], accuracy=91.8% at 3k frames
(b) LLama-VID [33], accuracy=55.0% at 10k frames
(c) VideoChat-Flash (ours), accuracy=99.1% at 10k frames
Figure 6: Results on the single-hop Needle-in-a-haystack evaluation with 10,000 frames.
Figure 7: Results on the multi-hop Needle-in-a-haystack with 10,000 frames.
@@
Fig. 7 shows that VideoChat-Flash beats baselines non-trivially in multi-hop NIAH. In evaluations, MLLMs need to trace along the chain of indicating images, locate the needle, and answer its question. Two metrics “CAP” and “’QA’ are used to denote the accuracy of finding the correct needle and the accuracy of answering the questions related to the needle as well as finding the needle successfully, respectively. As in the single-hop setting, we also sample 10,000 frames for testing. Specifically, VideoChat-Flash gives 31.3% and 25.4% in “CAP” and “QA” on average, higher than LongVA by around 8 points. Note that LongVA only works within 3000 frames. Compared with single-hop experiments, this shows VideoChat-Flash’s high-performing multimodal context in memorizing and understanding the quite long frame sequence.
@@ 4.3 Ablation & Analysis
We ablate key designs including hierarchical compression in VideoChat-Flash as follows.
Dynamic video sampling strategy.
As shown in Tab. 4, setting a lower bound of sampling frame number during training can significantly improve the fine-grained understanding in short videos. When the frame number increases from 8 to 64, the performance on MVBench is improved by 2.3%. Moreover, a varying sampling number for long videos is superior to a fixed number.
Training Testing MVBench VideoMME
Min Max Min Max Avg Overall
8 64 64 64 68.1 58.0
32 64 64 64 70.2 57.9
64 64 64 64 71.0 58.3
64 128 64 64 71.4 58.8
64 128 64 128 71.7 59.3
64 128 128 128 70.9 60.3
128 128 64 64 70.9 58.0
128 128 128 128 71.4 58.3
Table 4: Ablation of video sampling method.
Video encoders are efficient clip encoders.
Recent video MLLMs almost all use image encoders (e.g. SigLip) for video encoding. How image encoders and video ones perform in the context of video understanding is relatively under-explored. We empirically compare them in both short and long video evaluations and confirm that video encoder is still effective in encoding clips. As given in Tab. 5, we test the most popular image encoder, SigLIP [70], and the short video encoder, UMT [30], under different compressors. We find that even when the computational load is only a quarter of that of SigLIP, UMT still demonstrates remarkable performance and performs better on MVBench, demonstrating the merit of a video encoder in detailed temporal modeling.
Model FLOPs Latency MVBench VideoMME
(G) (ms) Avg Overall
w/ Spatial Downsampling, #tokens per frame=49
SigLIPSO400M@384 [70] 2679 79.7 60.2 53.6
UMT-L@224 [30] 596 11.8 61.5 53.1
w/ Resampler, #tokens per frame=49
SigLIPSO400M@384 [70] 2679 79.7 55.5 51.5
UMT-L@224 [30] 596 11.8 59.5 51.9
w/ Token Merging, #tokens per frame=49
SigLIPSO400M@384 [70] 2679 79.7 61.4 55.3
UMT-L@224 [30] 596 11.8 61.9 53.6
w/ Token Merging, #tokens per frame=16
SigLIPSO400M@384 [70] 2679 79.7 60.2 53.5
UMT-L@224 [30] 596 11.8 61.1 53.2
@@
Comparison of different visual compressors.
As shown in the Tab. 6, we found that the simple spatial downsampling performs better than the spatiotemporal resampler under various compression ratios. This indicates that it might be rather difficult to obtain a good compressor in a learnable manner with the currently relatively low amount of training data. The uneven downsampling, which utilizes the prior knowledge of inter-frame redundancy, can further enhance the effect of spatial downsampling. Meanwhile, directly merging similar tokens by leveraging the similarity among tokens is the most effective. When combined with the unfrozen visual encoder during training, a good compact spatiotemporal representation can be learned, and it can even outperform the results without compression when the compression ratio is low.
Compreesor MVBench MLVU VideoMME Avg
@@
#tokens per frame=729
Baseline (Uncompressed) 59.4 64 55.3 59.6
#tokens per frame=196
Spatial Downsampling 60.2 63.7 52.8 58.9
Uneven Downsampling 60.9 62.5 54.9 59.4
Spatio-temporal Resampler 59.5 61.9 51.9 57.8
Similar Token Merging 62.8 66.7 56.8 62.1
#tokens per frame=49
Spatial Downsampling 60.2 61.8 53.6 58.5
Uneven Downsampling 59.8 62.8 54.3 59.0
Spatio-temporal Resampler 55.5 58.1 51.1 54.9
Similar Token Merging 61.4 63.3 55.3 60.0
#tokens per frame=16
Spatial Downsampling 58.1 61.1 50.1 56.4
Uneven Downsampling 58.3 60.0 52.3 56.9
Spatio-temporal Resampler 51.4 54.7 47.7 51.3
Similar Token Merging 60.2 62.4 53.5 58.7
Table 6: Comparison of compressors.
@@
uniform/0.5 4 242.8 1.9 71.2 60.4
attention/0.5 4 242.8 1.9 70.7 60.8
uniform/0.5 18 295.2(-14%) 2.2 71.7 61.8
attention/0.5 18 295.2 2.2 72.1(+0.3) 61.7(+0.5)
attention/0.75,attention/0.25 4,18 245.8 1.9 71.4 60.9
uniform/0.75,attention/0.25 4,18 245.8(-28%) 1.9 72.0(+0.2) 61.1(-0.1)
Table 7: Effectiveness of visual dropout.
Multi-stage Short-to-Long Learning.
In Tab. 8, using a large amount of short footage for pre-training significantly improves the understanding of short videos as well as long ones. Incorporating both long videos and short videos in supervised fine-tuning (sft) of the model further enhances long video understanding while maintaining the ability to understand short videos. Additionally, adding the proposed long video data in Sec. 3.3 notably improves the performance in long video benchmark MLVU, at nearly no cost of performance drop in short video benchmark MVBench.
Settings MVBench MLVU VideoMME
Avg M-Avg Overall
Baseline (short video SFT) 61.5 61.6 53.1
$+$short video pretraining 65.8 63.1 54.3
$+$long video post-FT 67.0 64.5 55.5
$+$joint short & long video FT 67.0 64.5 55.5
$+$our long video SFT data 66.5 68.3 55.8
Table 8: Ablation of training strategies and data.

### v4

@@ 3.4 Multi-Hop Needle in A Video Haystack
Figure 4: An example of our Multi-Hop Needle in a Video Haystack. The right path (1, 2, 3) is for finding the needle while the wrong path (X1, X2, X3) is for distraction. MLLMs are asked to both find the needle (Q1) and answer its related question (Q2).
Previous works [68, 61] utilize the “Needle in a Video Haystack” (NIAH-Video) to evaluate the long video context understanding ability of models. Specifically, an image (commonly referred to as the “needle”) was inserted into a long video, and the model under test was then required to input the entire video and answer questions related to the needle. NIAH-Video assesses the model’s capability to retrieve information from long videos. However, it has several drawbacks. Firstly, it is difficult to prevent images and questions similar to the needle from appearing in the model’s training data, which leads to information leakage. Secondly, merely examining the model’s visual retrieval ability is insufficient and lacks discrimination for evaluating its long video context understanding ability (many models can achieve an accuracy rate over 99%). There is a need to further evaluate its reasoning ability regarding the content.
To address the above issues, we have designed a new evaluation task called “Multi-Hop Needle in a Haystack” (MH-NIAH-Video). As shown in Fig. 4, we insert a reasoning path composed of multiple images into the video haystack. Each image in this path has a randomly insertion position and corresponding textual clues to help find the next image. Given the starting point of the reasoning path, the model needs to follow this path to find the needle and answer questions related to it. What’s more, to prevent the model from skipping the step of finding the needle by relying on information leakage or memorizing the content of all images, we insert multiple wrong reasoning paths simultaneously while inserting the correct reasoning path. The model needs to find the correct needle (Q1) along the correct reasoning path based on the given starting point and then answer questions related to the needle (Q2). In a way, our multi-hop approach offers a much more robust evaluation of the long context understanding ability in Multimodal Large Language Models (MLLMs) compared to the previous NIAH-Video. In practice, all images are sourced from MS-COCO [32], making use of its human-annotated captions and question-answer pairs. It should be noted that even if the model can perfectly remember the content of MS-COCO, it will not be of much help in finding the needle, which significantly reduces the likelihood of successful “cheating”.
@@ 4 Experiments
Model Size Avg tokens MVBench PerceptionTest LongVideoBench MLVU VideoMME (w/o & w sub.) LVBench Charades-STA AuroraCap per frame Avg Val Val M-Avg Overall Long Avg mIoU Avg Avg. Duration 16s 23s 473s 651s 1010s 2386s 4101s 30s 28s Proprietary Models GPT-4V [39] - - 43.7 - 59.1 49.2 59.9/63.3 53.5/56.9 - - - GPT-4o [40] - - 64.6 - 66.7 64.6 71.9/77.2 65.3/72.1 30.8 35.7 - Gemini-1.5-Pro [44] - - 60.5 - 64.0 - 75.0/81.3 67.4/77.4 33.1 - 41.7 Small Size MLLMs Qwen2-VL [52] 2B 1924 63.2 - - - 55.6/60.4 - - - - InternVL2.5 [12] 2B 256 68.8 - 46.0 61.4 51.9/54.1 - - - - VideoChat-Flash @448 2B 16 70.0 70.5 58.3 65.7 57.0/63.9 44.9/54.0 42.9 45.2 - Open-Source MLLMs VideoChat2-HD [29] 7B 72 62.3 - - 47.9 45.3/55.7 39.8/53.9 - 3.4 - InternVideo2-HD [56] 7B 72 67.2 63.4 - - 49.4/ - - - - - LLaVA-OneVision [25] 7B 196 56.7 57.1 56.3 64.7 58.2/61.5 - - 13.5 37.5 LLaVA-OneVision [25] 72B 196 59.4 66.9 61.3 68.0 66.2/69.5 - - - - LLaVA-Video [71] 7B 676 58.6 67.9 58.2 70.8 63.3/69.7 - - - 39.0 VITA1.5 [17] 7B 256 56.8 - - - 56.8/59.5 - - - - - InternVL2 [13] 8B 256 65.8 - 54.6 64.0 54.0/56.9 - - - 37.7 InternVL2 [13] 76B 256 69.6 - 61.1 69.9 61.2/62.8 - - - - InternVL2.5 [12] 8B 256 72.0 - 60.0 68.9 64.2/66.9 - - - - Qwen2-VL [52] 7B 1924 67.0 66.9 - - 63.3/69.0 - - - 41.6 Qwen2.5-VL [3] 7B 1924 69.6 - 56.0 70.2 65.1/71.6 - 45.3 43.6 - Open-Source Long Video MLLMs LLaMA-VID [30] 7B 2 41.9 44.6 - 33.2 25.9/ - - 23.9 - 30.9 LongVU [47] 7B 64 66.9 - - 65.4 - /60.6 - /59.5 - - - LongVA [68] 7B 144 - - - 56.3 52.6/54.3 46.2/47.6 - - 34.5 LongVILA [61] 7B 196 67.1 58.1 57.1 - 60.1/65.6 47.0/52.1 - - - Kangaroo [34] 8B 256 61.0 - 54.8 61.0 56.0 / 57.6 46.7 / 49.3 39.4 - - VideoChat-Flash @224 7B 16 73.2 75.6 64.2 74.5 64.0/69.4 53.6/61.9 47.2 48.4 - VideoChat-Flash @448 7B 16 74.0 76.2 64.7 74.7 65.3/69.7 55.4/63.3 48.2 48.0 42.9
Table 1: Results on comprehensive video-linguistic benchmarks
@@
We employ UMT-L [28], token merging with MLP, and Qwen2-7B as visual encoder, connector, and LLM, respectively. When processing a long video, we divide it into shorter clips, each consisting of 4 frames. Each clip is compressed into 64 tokens, meaning that, on average, each frame is represented by 16 tokens. Regarding video-level compression, while it presents some challenges in compatibility with training acceleration strategies such as sequence parallelism, we only employ it during inference. In most of the ablations, we use only one-fourth of the full dataset. See Appendix for details.
@@ 4.1 General Video Understanding Evaluation
We evaluate our model on six general video understanding benchmarks in question-answering format, including two short video benchmarks: MVBench [29] and Perception Test [41], and three long video benchmarks: LongVideoBench [59], MLVU [73] and LVBench [53], and a comprehensive benchmark, VideoMME [16], covering videos ranging from minute-level to hour-level durations. We further evaluate the temporal grounding and video caption tasks, using the Charades-STA [18] and AuroraCap [21].
Leading performance.
As depicted in Tab. 1, our VideoChat-Flash achieves the best results on diverse VideoQA benchmarks within the 2B and 7B size category, significantly outperforming other approaches. Remarkably, its performance even eclipses that of models with substantially larger scales, such as InternVL2-76B, as well as proprietary models like GPT-4o and Gemini-1.5-Pro. Even when merely supplying timestamp information via a text prompt, our model has achieved remarkable performance in temporal grounding. Meanwhile, it also significantly outperforms other models in the video captioning task, even surpassing the proprietary GPT-4o and Gemini-1.5 Pro. This demonstrates the effectiveness of the comprehensive design of our model, data, and training strategies.
(a) LongVA [68], accuracy=91.8% at 3k frames
(b) LLama-VID [30], accuracy=55.0% at 10k frames
(c) VideoChat-Flash (ours), accuracy=99.1% at 10k frames
Figure 5: Results on the “Single-Hop Needle-in-A-Video-Haystack” evaluation with 10,000 frames.
4.2 Long Video Context Evaluation
@@
As shown in Fig. 5, we follow the protocols in LongVA [68] for Single-Hop NIAH, we source a long video and sample frames uniformly from it. Then we add needles (indicating images) into the sampled image sequence at different positions. MLLMs are fed with this long image sequence and answer the corresponding questions to the indicating images. We evaluate all models over 10,000 frames. Note our VideoChat-Flash delivers a 99.1% success rate in accurately retrieving the correct indicating image and answering the related question even across 10,000 frames. In comparison, LongVA gives a decent result close to 92% within 3000 frames while LLama-VID only achieves 55% accuracy. It demonstrates VideoChat-Flash’s state-of-the-art performance in long multimodal context modeling.
Figure 6: Results on the “Multi-Hop Needle-in-A-Video-Haystack with 10,000 frames.
@@
In this evaluations, MLLMs need to trace along the chain of indicating images, locate the needle, and answer its question. Two metrics “CAP” and “QA” are used to denote the accuracy of finding the correct needle and the accuracy of answering the questions related to the needle as well as finding the needle successfully, respectively. As shown in Fig. 6, our VideoChat-Flash still beats all baselines. Specifically, VideoChat-Flash gives 31.3% and 25.4% in “CAP” and “QA” on average, higher than LongVA by around 8 points. It can be seen that compared with the single-hop NIAH, the multi-Hop NIAH presents a much more difficult challenge, which can better reflect the real gap between the capabilities of different models.
@@ 4.3 Ablation & Analysis
Effect of various designs.
As shown in Tab. 2, we have conducted comprehensive ablation studies on each design. In terms of the model, it can be observed that HiCo significantly reduces the computational load (from 196 to 16 tokens per frame) while barely compromising the performance. Meanwhile, duration-based sampling and timestamp prompts play crucial roles in enhancing the performance. The further leap in performance mainly stems from the training strategy in short-to-long learning and a better mixture of training data.
Settings MVB MLVU VMME Charades
Avg M-Avg Overall mIoU
Baseline 60.2 63.7 52.8
+ HiCo 61.1 60.6 53.2 -
+ short video pretraining 66.5 62.4 53.9 -
+ duration-based sampling 67.0 64.5 55.5 -
+ LongVid data 66.5 68.3 55.8 -
+ Joint short & long sft 73.2 74.5 64.0 48.4
+ High-res post ft 74.0 74.7 65.3 48.0
- timestamp prompt 73.4 73.2 63.4 44.2
Table 2: Effect of various designs on data , model , and resolution . The baseline employs SigLiP-so400M [65] as the vision encoder and Spatial donwsampling (196 tokens per frame) as the connector. It adopts a two-stage training strateay with image and short video following LLaVA [33].
Duration-based Sampling.
As shown in Fig. 7, A relatively large $T_{\text{min}}$ (64) enables the model to better learn to model the fine actions and rapid movements in short videos during training, thereby enhancing the performance of short video understanding. Increasing $T_{\text{max}}$ from 64 to 256 leads to a stable improvement in the understanding performance of both short and long videos. This indicates that more sampled frames can extract more accurate information from our long video data. When $T_{\text{max}}$ reaches 512, there is a slight decline in the performance of short videos. Overall, it achieves a balance between the performance of short and long videos.
(a) Effect of $T_{\text{min}}$ ($T_{\text{max}}$=128)
(b) Effect of $T_{\text{max}}$ ($T_{\text{min}}$=64)
Figure 7: Ablation of Duration-based Sampling
Video encoders are efficient clip compressors.
As shown in Tab. 3, we have tested the most popular image encoder, SigLIP [65], and the short video encoder, UMT [28], for encoding clips with heavy compression. We found that even when the computational cost is significantly lower, UMT can still achieve better performance on the short video task MVBench. Moreover, as the size of the training data increases from 2 million to 8 million, UMT outperforms SigLIP distinctly across various benchmarks. We believe that this is attributed to the spatio-temporal attention employed by UMT, which can aggregate the key information from different frames within a clip, thus enabling the learning of more compact compression features.
Visual Encoder FLOPs Latency MVBench MLVU VideoMME
(G) (ms) Avg M-Avg Overall
#tokens per frame=16, training data size=2M
SigLIPSO400M@384 2679 79.7 60.2 62.0 53.5
UMT-L@224 596 11.8 61.1(+0.9) 60.0(-2.0) 53.2(-0.3)
#tokens per frame=16, training data size=8M
SigLIPSO400M@384 2679 79.7 71.2 70.8 62.4
UMT-L@224 596 11.8 73.5(+2.3) 73.7(+2.9) 62.7(+0.3)
@@
Different connectors and compression ratio.
As shown in the Tab. 4, we consider three different numbers of tokens per frame after compression (16, 49, 196) and four popular token compression strategies: spatial downsampling [70, 13], uneven downsampling [57], spatio-temporal resampler [29, 50], and similar token merging [6, 58] (more details can be found in the Appendix). It can be seen that compared with other methods, the parameter-free similar token merging operation can achieve a remarkably low compression ratio and even obtain better performance than without compression. Even in the extreme case of a 2% compression ratio, it can still maintain most of the performance.
Connector MVBench MLVU VideoMME Avg
@@
#tokens per frame=729, compression ratio=100%
MLP (Uncompressed) 59.4 64 55.3 59.6
#tokens per frame=196, compression ratio=27%
Spatial Downsampling 60.2 63.7 52.8 58.9(-0.7)
Uneven Downsampling 60.9 62.5 54.9 59.4(-0.2)
Spatio-temporal Resampler 59.5 61.9 51.9 57.8(-1.8)
Similar Token Merging 62.8 66.7 56.8 62.1(+2.5)
#tokens per frame=49, compression ratio=7%
Spatial Downsampling 60.2 61.8 53.6 58.5(-1.1)
Uneven Downsampling 59.8 62.8 54.3 59.0(-0.6)
Spatio-temporal Resampler 55.5 58.1 51.1 54.9(-4.7)
Similar Token Merging 61.4 63.3 55.3 60.0(+0.4)
#tokens per frame=16, compression ratio=2%
Spatial Downsampling 58.1 61.1 50.1 56.4(-3.2)
Uneven Downsampling 58.3 60.0 52.3 56.9(-2.7)
Spatio-temporal Resampler 51.4 54.7 47.7 51.3(-8.3)
Similar Token Merging 60.2 62.4 53.5 58.7(-0.9)
Table 4: Comparison of connectors.
@@
Uni./0.5 4 242.8 1.9 71.2 60.4
Attn./0.5 4 242.8 1.9 70.7 60.8
Uni./0.5 18 295.2 2.2 71.7 61.8
Attn./0.5 18 295.2 2.2 72.1(+0.3) 61.7(+0.5)
Attn./0.75,Attn./0.25 4,18 245.8 1.9 71.4 60.9
Uni./0.75,Attn./0.25 4,18 245.8 1.9 72.0(+0.2) 61.1(-0.1)
Table 5: Effectiveness of visual dropout. The Qwen2-7B we used has a total of 28 layers. ”Uni.” and ”Attn.” represent uniform drop and attention select respectively.
Input Model Avg tokens FLOPs Memory(G)
frames per frame (T) Train Infer
64 LongVILA [61] 196 224.8 15.4 16.7
LongVA [68] 144 155.9 12.3 16.3
VideoChat-Flash 16 14.8 4.8 15.4
256 LongVILA [61] 196 1467.5 50.1 21.0
LongVA [68] 144 930.4 37.8 19.5
VideoChat-Flash 16 63.0 7.6 15.7
1000 LongVILA [61] 196 14336.9 oom 37.7
LongVA [68] 144 8278.9 oom 31.8
VideoChat-Flash 16 303.3 18.6 17.1
10000 LongVILA [61] 196 1184250.0 oom oom
LongVA [68] 144 644632.0 oom oom
VideoChat-Flash 16 9969.5 oom 33.6
Table 6: Comparison of FLOPs and Cuda memory. The FLOPs and inference memory is estimated using one NVIDIA A100-80G GPU with one sample, and the training is estimated using 32 NVIDIA A100-80G GPUs with DeepSpeed ZeRO-3 [42]. We assume that the visual features have been extracted and stored in advance, so we only consider the FLOPs and memory of the LLM.
Model efficiency.
As in Tab. 6, even when processing short videos, the compute load of our model is only one-tenth that of previous models. Meanwhile, as the number of input frames increases, the difference becomes more and more pronounced. Only our model can complete the inference on 10,000 frames on a single A100-80G. Concretely, VideoChat-Flash’s compute load is two orders of magnitude lower than that of LongVILA [61] (9,969.5 vs. 1,184,250.0 TFLOPs).